A research institute conducted clinical trials of a method designed to increase the probability of conceiving a boy. Among 147 babies born to parents using the​ method, 118 were boys. Identify the margin of error and the​ 95% confidence interval for these clinical trials.

The margin of error is____. ​(Round to three decimal places as​ needed.)

Construct the​ 95% confidence interval.

_____< p < _____

(Round to three decimal places as​ needed.)
Answer #1

X=   118  
n=   147  
p-hat = X/n=   0.803    =118/147

critical value, z(a/2) = z(0.05/2) = 1.960

ME = z(a/2)*sqrt(p*(1-p)/n)
ME = 1.96*SQRT(0.803*(1-0.803)/147) = 0.064

CI = p +- ME
lower = 0.803-0.0643 = 0.739
upper = 0.803+0.0643 = 0.867